OpenStudy (3psilon):

If one root of a quadratic equation is 2 - i , what is the equation?

hartnn (hartnn):

other root = 2+i

hartnn (hartnn):

(x-(2-i))(x-(2+i)) can u solve further >?

OpenStudy (3psilon):

How do we know that's the other root?

hartnn (hartnn):

COMPLEX ROOTS ALWAYS OCCUR IN PAIRS if a+ib is one root, a-ib is other root

OpenStudy (3psilon):

Okay I'm remembering that now! But how do we formulate the equation?

hartnn (hartnn):

if a is the root, x-a is the factor

OpenStudy (3psilon):

Can you explain further please

hartnn (hartnn):

since 2-i is a root, x-(2-i) is a factor since 2+i is a root, x-(2+i) is a factor since its a quadratic equation, it only has 2 factors which are (x-(2-i))(x-(2+i)) now just multiply them out

hartnn (hartnn):

or use (a-b)(a+b) = a^2-b^2

hartnn (hartnn):

[(x-2) + i ][(x-2)-i] = (x-2)^2 - i^2 getting this ?

OpenStudy (3psilon):

I'm piecing it together give me a sec

OpenStudy (3psilon):

I got x^2 - 4x + 5 = 0! Wow that's amazing! Thank you so much! Explains it much better than Dr. Chung's book

hartnn (hartnn):

lol, welcome ^_^